OpenStudy (anonymous):

Solve for x: |2x + 6| − 4 = 20

OpenStudy (jhannybean):

isolate |2x+6| to one side by adding +4 to both sides. \[|2x+6| = 24\]

pooja195 (pooja195):

Add 4 to both sides. |2x+6|−4+4=20+4 |2x+6|=24 Solve Absolute Value. |2x+6|=24 We know either 2x+6=24 or 2x+6=−24

OpenStudy (anonymous):

isolate |2x+6| to one side by adding +4 to both sides. what do you think is anwser

OpenStudy (jhannybean):

Now the absolute value means that x has 2 answers, -(2x+6) =24 or 2x+6 =24